# Camouflaged Object Tracking: A Benchmark

### v3

@@ I Introduction
In recent years, the field of visual tracking has made significant advances, largely due to the widespread use of large-scale training datasets. The richness of these datasets has facilitated the development of high-precision and robust tracking algorithms [1, 2, 3, 4, 5, 6, 7, 8]. Although substantial progress has been made in general object tracking, research into more specialized and challenging scenarios remains limited. Among these, tracking camouflaged objects is particularly difficult because these targets seamlessly blend into their surroundings, making them hard to detect with the naked eye and significantly increasing tracking complexity [9, 10]. While existing tracking methods have achieved impressive results on regular objects through various techniques such as high-level semantics [11], block-based correlation filters [12], spatio-temporal feature learning [13], and long-term tracking frameworks [14], their performance degrades significantly when dealing with camouflaged targets due to the inherent visual ambiguity between the target and background.
@@
Distinguishing camouflaged objects from their backgrounds in these applications is crucial, but it involves significant challenges like background clutter, boundary blurring, texture mimicry, color similarity, varying lighting conditions, and occlusion [25, 9, 10]. Addressing these challenges through innovative approaches and advanced technologies can lead to significant advancements in applications involving camouflaged objects [26, 27, 28, 29, 30]. In recent years, numerous computer vision techniques have been developed to address the challenges posed by camouflaged objects [31, 30]. One prominent direction is Camouflaged Object Detection (COD), which identifies objects blending into their surroundings due to similar textures, shapes, or colors. COD locates these hidden objects within images or videos [32, 33]. Camouflaged Object Segmentation (COS) takes this further by not only detecting but also accurately separating these objects from their backgrounds. COS involves classifying each pixel to determine if it belongs to the object or the background, offering detailed information about the object’s shape and edges [34]. The ongoing research in COD and COS continue to push the boundaries, promising further innovations and applications that tackle real-world challenges related to camouflaged objects.
@@
Advancements in COD and COS technologies have been significantly propelled by the integration of deep learning algorithms. Methods such as MirrorNet [34], PFNet [25], and S-MGL [35] have greatly improved detection and segmentation accuracy by leveraging vast datasets to learn and generalize the subtle differences between camouflaged objects and their backgrounds. To support the development and evaluation of these algorithms, large-scale benchmark datasets like NC4K [30] and COD10K [32] have been established, providing crucial resources for advancing the field. Despite these advancements, Camouflaged Object Tracking (COT), which focuses on maintaining consistent identification of camouflaged objects throughout a video while accounting for changes in position, scale, and appearance, remains under-explored. This task requires not only the detection of camouflaged targets but also the continuous tracking of their location and status across video streams. To achieve this, more refined visual feature extraction and sophisticated tracking algorithms are necessary to prevent target loss or misalignment. Additionally, the inherent uncertainty and instability of target motion and the presence of coexisting objects further complicate the tracking process [36, 37, 38]. These challenges highlight the need for substantial innovation and effort to effectively address the complexities of COT in real-world scenarios. However, there is a critical gap in the field: to our knowledge, no specialized dataset exists for camouflaged object tracking. This absence significantly limits the development and rigorous evaluation of COT algorithms, as existing datasets may not adequately represent the unique challenges posed by camouflaged objects. Without a dedicated dataset, researchers face difficulties in testing their models against the specific nuances of camouflaged objects.
@@
In this work, we take the first step in addressing the challenging task of tracking camouflaged objects by introducing the Camouflaged Object Tracking Dataset (COTD). COTD features a diverse range of camouflaged objects across various environments, specifically designed to push the boundaries of camouflaged object tracking algorithms. The dataset includes 200 sequences totaling around 80,000 frames, with each sequence meticulously annotated with detailed bounding boxes and relevant attributes for comprehensive performance evaluation and analysis. To the best of our knowledge, COTD is the first benchmark dataset specifically dedicated to the task of tracking camouflaged objects. Fig. 1 demonstrates the differences between generic object tracking and camouflaged object tracking. Compared with generic object tracking, where targets are more distinguishable from their backgrounds, tracking camouflaged objects is notably more challenging as these objects blend seamlessly with their surroundings. This blending creates ambiguous visual cues that significantly complicate the tracking process. Our extensive dataset captures a broad spectrum of real-world scenarios involving camouflaged objects, laying a solid foundation for developing and evaluating new tracking algorithms that adapt to these conditions. We also conducted extensive evaluations of 20 state-of-the-art tracking algorithms and found that current methods often struggle with the complexities introduced by camouflaged objects, indicating the need for substantial improvements. By releasing COTD, we aim to advance tracking technology to address the unique challenges posed by camouflaged targets and ultimately enhance the accuracy and reliability of visual tracking systems across various applications. The main contributions of this paper are as follows:
@@ II-A Visual Tracking Algorithms
DCF-based trackers treat the visual tracking problem as an online regression task, leveraging the Parseval theorem and Fast Fourier Transform (FFT) for efficient computation. The MOSSE (Minimum Output Sum of Squared Error) filter [39] is an early representative in this domain. Since its introduction, significant advancements have been made. For example, [40, 41] introduced additional scale filters to address variations in object scale, while [42] and [43] employed regularization techniques to enhance tracker robustness. Furthermore, extensions of the DCF framework, such as translation-invariant techniques, continue to evolve [44]. However, due to the limited representational capacity of handcrafted features, DCF-based trackers still face challenges in robustness under complex conditions.
To overcome the limitations of DCF-based trackers, deep learning-based trackers have garnered significant attention in recent years. These trackers utilize neural networks to automatically learn features, substantially improving tracking performance. SiamFC [45] pioneered this field by applying Siamese networks [46] to visual tracking. SiamFC learns a nonlinear similarity function through offline training and performs correlation operations on candidate regions and the initial object state to generate response maps for the search area. Building on this, the SiamRPN algorithm introduced a candidate region generation network structure with classification and regression branches to enhance tracking speed. Subsequent research, including SiamRPN++ [47], SiamR-CNN [48], SiamCAR [49], and Ocean [50], has enhanced the accuracy and robustness of Siamese trackers through techniques like attention mechanisms in SiamCAR, backbone optimization in SiamRPN++, ResNet-50 feature enhancement in Ocean, and two-stage detection in SiamR-CNN.
Recently, the emergence of Vision Transformers (ViTs) [51] has provided new directions for the development of visual tracking technologies. Vision Transformers demonstrate potential in simplifying and unifying tracking frameworks [51, 52, 53, 54]. For instance, OStrack [54] integrates feature extraction and relational modeling, and introduces an early elimination module to achieve an efficient single-stream tracking framework. DropTrack [53] improves temporal correspondence learning by adaptively performing spatial attention dropout during frame reconstruction. HIPTrack [55] introduces a history prompting network that utilizes refined historical foreground masks and object visual features, providing comprehensive and precise guidance for trackers.
@@ II-B Visual Tracking Benchmarks
Over the past decade, numerous benchmarks have been proposed for testing and evaluating single-object tracking performance. The advancement of visual tracking technology relies on accurate measurement and evaluation of tracking methods, with benchmark datasets serving as a critical evaluation standard. Currently, there are two main types of benchmark datasets: general benchmarks and specific benchmarks [56].
General benchmarks: General benchmark datasets are primarily used to assess the performance of tracking algorithms in typical scenarios. OTB [57, 58] and VOT [59, 60, 61, 62, 63, 64] are among the initial platforms for unified testing data and evaluation standards for general visual tracking. NFS [65] focuses on high-frame-rate videos. Many large-scale benchmarks have been introduced for training deep learning-based trackers. For example, TrackingNet [66] aims to provide resources for training and evaluating general object tracking algorithms, containing over 30,000 sequences. GOT-10K [67] includes over 10,000 videos from the WordNet semantic hierarchy to develop the generalization capability of trackers. LaSOT [68] encompasses 1,400 long-term video sequences, with an additional 150 sequences and a new evaluation scheme.
Specific benchmarks: Specific benchmark datasets aim to evaluate tracking performance for particular types of objects. For example, UAV123 [69] focuses on low-altitude drone object tracking, containing 123 sequences captured by drones. CDTB [70] and PTB [71] are designed to evaluate tracking performance in RGB-D videos, where D represents depth images. TOTB [56] specializes in tracking transparent objects, collecting 225 videos from 15 categories of transparent objects. Recently, TSFMO [72] has been introduced for tracking small and fast-moving objects, while TRO [73] focuses on tracking reflected objects.Additionally, BihOT [74] contains 49 videos and 42K frames, focusing on hyperspectral camouflage. UW-COT220 [75] focuses on underwater camouflage
Despite the availability of existing benchmarks, they mainly focus on objects that do not intentionally blend into their surroundings. Camouflaged objects, however, present a unique challenge that these benchmarks do not adequately address. To fill this critical gap, we propose the Camouflaged Object Tracking Dataset (COTD) in this work.
@@ II-C Dealing with camouflaged Objects in Vision
Camouflaged objects are difficult to distinguish from their backgrounds because they blend into their surroundings. Unlike regular objects, they have similar textures, colors, and patterns as their environment, creating visual ambiguity. Recently, many computer vision techniques have been developed to tackle these challenges. For instance, Fan et al. [32] introduced the task of Camouflaged Object Detection (COD) and established a dataset COD10K with 10,000 images, along with a simple yet effective framework SINet for efficient and precise detection of camouflaged objects. Liu et al. [33] proposed a scale-adaptive correlation filter algorithm that integrates FHOG and LBP features for Camouflaged Object Segmentation (COS). Yue et al. [76] and Chen et al. [77] have further refined segmentation accuracy by integrating multi-scale feature fusion techniques. Cheng et al. [26] developed a Video Camouflaged Object Detection (VCOD) framework that captures inter-frame motion and uses a spatio-temporal transformer for consistency, along with the MoCA-Mask dataset providing detailed pixel-level annotations. Additional research by Wang et al. [78] have demonstrated the potential of leveraging discriminative feature learning for enhancing detection robustness in complex environments. Despite these advancements, Camouflaged Object Tracking (COT), which focuses on consistently identifying camouflaged objects in videos while adapting to changes in position, scale, and appearance, remains under-explored. The absence of publicly available datasets for camouflaged object tracking has further constrained the progress of research and algorithm development in this field. To bridge this gap, we introduce the Camouflaged Object Tracking Dataset (COTD), specifically designed to support the development and evaluation of algorithms tailored for tracking camouflaged objects.
@@ III-A Data Collection
Figure 3: This chart depicts the distribution of sequence counts for 20 categories. Each category is represented by a bubble, where the size of the bubble corresponds to the number of sequences in that category. Larger bubbles indicate categories with higher sequence counts, visually highlighting the differences in quantity among the categories.
@@
In the video selection process, we apply stringent criteria to ensure the high quality and practicality of the data. Specifically, videos are evaluated and selected based on three main criteria: 1) the diversity and complexity of the environment, 2) the frequency with which targets cross frame boundaries, and 3) the overall length and continuity of the videos. We prioritize videos that present additional challenges, such as a high degree of similarity between targets and backgrounds. Due to resource constraints in collecting, processing, and annotating video data, we select 200 sequences for the benchmark test. This ensures that each video is of high quality with comprehensive annotations, adhering to our available time, budget, and manpower constraints. The COTD dataset comprises 20 categories, namely: bird, bug, butterfly, cat, crab, dog, fish, frog, grasshopper, leopard, lizard, mantis, octopus, person, phasmatodea, seahorse, rabbit, shrimp, snake, and spider. To visually represent the composition of the dataset, we create a chart illustrating the number of sequences per category. Additionally, we generate a series of graphs depicting the maximum, minimum, and average number of frames for each category. These visual aids not only enhance the understanding of the dataset’s diversity but also highlight the challenges in tracking camouflaged objects across different scenarios. Fig. 3 presents the distribution of sequences across the 20 categories, while Fig. 4 provides a detailed view of the frame characteristics for each category, such as the range and average number of frames. These figures are crucial in demonstrating the dataset’s comprehensiveness and its potential as a robust benchmark for evaluating the performance of camouflaged object tracking algorithms. Additionally, to highlight the challenges associated with tracking camouflaged objects, we evaluate seven state-of-the-art trackers on the COTD dataset and compare them with two popular generic object tracking benchmarks. The AUC results are presented in Table I. Notably, the AUC of these trackers on the TrackingNet benchmark is more than 15.0% higher than on the COTD dataset, underscoring the significant challenges posed by tracking camouflaged objects.
Figure 4: The radar chart compares video frame counts (minimum, average, maximum) across 20 categories. Blue, orange, and green lines represent the distributions of minimum, average, and maximum frame counts respectively for each category, illustrating differences between them.
@@
(a) Tight bounding box on the first
frame.
(b) The fish is partially occluded by the hand.
(c) The fish is fully occluded by the hand.
(d) The fish partially passed the hand.
Figure 5: Illustration of our annotation rules using fish-19, where the fish performs a jump and lands on the hand. In (a), the bounding box is kept as tight as possible around the fish before it jumps. In (b) and (d), the fish is partially obscured by the person’s hand. We provide tight bounding boxes that closely follow the visible parts of the fish. In (c), the fish is fully occluded by the hand, and thus no bounding box is provided due to the Full Occlusion (FO) scenario.
@@ III-B Data Annotation
We follow the principles in [66] for sequence annotation. Each frame undergoes meticulous annotation by expert annotators, comprising students specializing in visual tracking research. The annotation process involves drawing axis-aligned bounding boxes that represent the most compact boundary of the target object, with careful documentation of object visibility, occlusion, and absence states. Examples of bounding box annotations in the COTD are shown in Fig. 2.
To ensure annotation quality and consistency, we implement a rigorous two-stage verification process. In the primary stage, annotators conduct frame-by-frame labeling, ensuring precise object identification even in complex scenes. This is followed by a secondary verification where another annotator reviews all annotations, with any discrepancies resolved through discussion and consensus.
The details of our annotation rules are as follows.
Rule 1: All visible components of the designated object must be encompassed within the bounding box, which should be drawn as tightly as possible around the object.
Rule 2: In instances where the target object is entirely occluded or exits the frame, no bounding box should be provided for that particular frame.
Rule 3: A valid annotated bounding box must be present in the initial frame of each video sequence to delineate the target object.
Rule 4: It is imperative to document the entire gradual transition of the target object when it undergoes occlusion, disappearance, or re-emergence within the scene.
As illustrated in Fig. 5, we adhere to the rules by drawing bounding boxes that are tightly fitted within the boundaries of the target object, avoiding overly large sizes to ensure precision across numerous frames. Additionally, we focus on accurately annotating the processes of the target object’s disappearance and re-emergence (either FO or OV), which is crucial for training trackers to handle occlusions and the recapture of targets. This meticulous annotation helps create specific test scenarios, such as long-term occlusions, which are very useful for evaluating the recapture capabilities of tracking algorithms.
@@ III-C Attributes
Figure 6: Distribution of sequences with each attribute in COTD.
Figure 7: The overall framework of HIPTrack-MLS, which is inherited from HIPTrack. The primary difference is in the encoder component of the feature extraction network. The original one is replaced by the proposed Multi-Level Feature Sensing Encoder.
Figure 8: Structure of the Multi-Level Feature Sensing Encoder.
Following the prevalent tracking benchmark methodologies [94], [67], and [58], we meticulously annotate each video sequence with detailed attributes to enable an in-depth analysis of tracking algorithms’ performance. The proposed COTD dataset features 12 attributes, with each sequence including Background Clutter (BC) to highlight the dataset’s camouflage characteristics. In addition, we annotate the following 11 attributes: (1) Illumination Variation (IV), (2) Scale Variation (SV), (3) Deformation (DEF), (4) Motion Blur (MB), (5) Fast Motion (FM), which is marked when the object center moves by at least 20% of its own size between consecutive frames, (6) Out-of-View (OV), (7) Low Resolution (LR), marked when the object area is smaller than 900 pixels, (8) Partial Occlusion (POC), (9) Rotation (ROT), (10) Full Occlusion (FOC), and (11) Aspect Ratio Change (ARC), marked when the object’s bounding box aspect ratio falls outside the range [0.5, 2]. Fig.6 shows the distribution of these attributes on COTD. As can be seen, the most common challenge in COTD is Motion Blur. In addition, the Deformation and Partial Occlusion also present frequently in COTD.
TABLE III: Co-occurrence distributions of the attributes in COTD. The diagonal demonstrates the distribution of the entire dataset, and each row or column represents the distribution of the corresponding attribute subset.
Attribute IV SV DEF MB FM OV LR POC ROT FOC ARC
IV 5 0 1 1 0 0 0 2 0 0 0
SV 0 29 14 12 7 8 3 17 2 1 4
DEF 1 14 59 32 9 16 5 27 13 3 8
MB 1 12 32 72 12 19 10 25 8 4 6
FM 0 7 9 12 14 5 2 8 3 2 1
OV 0 8 16 19 5 39 3 12 4 2 2
LR 0 3 5 10 2 3 21 6 0 0 0
POC 2 17 27 25 8 12 6 58 11 4 6
ROT 0 2 13 8 3 4 0 11 29 1 3
FOC 0 1 3 4 2 2 0 4 1 4 0
ARC 0 4 8 6 1 2 0 6 3 0 9
Based on the 11 attributes, the co-occurrence matrix of the COTD dataset is provided in Table III. The horizontal and vertical axes correspond to each attribute, and the numbers in the table refer to the number of sequences with both horizontal and vertical attributes. For example, the number of sequences with both MB and DEF is 32. By comprehensively considering these attributes, we can conduct a thorough evaluation of tracking algorithms under various challenging conditions.

### v4

@@ I Introduction
In recent years, the field of visual tracking has made significant advances, largely due to the widespread use of large-scale training datasets. The richness of these datasets has facilitated the development of high-precision and robust tracking algorithms [1, 2, 3, 4, 5, 6, 7, 8]. Although substantial progress has been made in general object tracking, research into more specialized and challenging scenarios remains limited. Among these, tracking camouflaged objects is particularly difficult because these targets seamlessly blend into their surroundings, making them hard to detect with the naked eye and significantly increasing tracking complexity [9, 10].
@@
Distinguishing camouflaged objects from backgrounds is crucial but challenging due to factors like background clutter, boundary blurring, texture mimicry, color similarity, varying lighting, and occlusion [21, 9, 10]. Addressing these challenges can significantly advance applications involving camouflaged objects [22, 23, 24, 25, 26]. Recent computer vision techniques, particularly in Camouflaged Object Detection (COD) and Camouflaged Object Segmentation (COS), have made notable progress. COD identifies objects blending into their surroundings, while COS further separates them from backgrounds by classifying each pixel [27, 28, 29]. Ongoing research in these areas continues to push boundaries, promising innovations to tackle real-world challenges related to camouflaged objects.
@@
Deep learning has significantly advanced COD and COS, with methods like MirrorNet [29], PFNet [21], and S-MGL [30] improving accuracy by learning from large datasets. Benchmark datasets such as NC4K [26] and COD10K [27] have been crucial for this progress. However, Camouflaged Object Tracking (COT), which involves consistent identification of camouflaged objects in videos, remains under-explored. COT requires refined feature extraction and sophisticated algorithms to handle changes in position, scale, and appearance, as well as to manage uncertainties and coexisting objects [31, 32, 33]. A critical gap exists: there is no specialized dataset for COT. This absence limits the development and evaluation of COT algorithms, as existing datasets may not capture the unique challenges of camouflaged objects. Without a dedicated dataset, researchers struggle to test their models against these specific nuances.
@@
In this work, we take the first step in addressing the challenging task of tracking camouflaged objects by introducing the Camouflaged Object Tracking Dataset (COTD). COTD features a diverse range of camouflaged objects across various environments, specifically designed to push the boundaries of camouflaged object tracking algorithms. The dataset includes 200 sequences totaling around 80,000 frames, with each sequence meticulously annotated with detailed bounding boxes and relevant attributes for comprehensive performance evaluation and analysis. Fig. 1 demonstrates the differences between generic object tracking and camouflaged object tracking. Compared with generic object tracking, where targets are more distinguishable from their backgrounds, tracking camouflaged objects is notably more challenging as these objects blend seamlessly with their surroundings. This blending creates ambiguous visual cues that significantly complicate the tracking process. Our extensive dataset captures a broad spectrum of real-world scenarios involving camouflaged objects, laying a solid foundation for developing and evaluating new tracking algorithms that adapt to these conditions. We also conducted extensive evaluations of 20 state-of-the-art tracking algorithms and found that current methods often struggle with the complexities introduced by camouflaged objects, indicating the need for substantial improvements. By releasing COTD, we aim to advance tracking technology to address the unique challenges posed by camouflaged targets and ultimately enhance the accuracy and reliability of visual tracking systems across various applications. The main contributions of this paper are as follows:
@@ II-A Visual Tracking Algorithms
DCF-based trackers, like the early MOSSE filter [34], treat visual tracking as an online regression problem. Enhancements such as scale filters and regularization techniques have been introduced to improve robustness [35, 36, 37, 38, 39]. However, these trackers still struggle with complex conditions due to their reliance on handcrafted features. To overcome these challenges, deep learning-based trackers have become popular. A key innovation was SiamFC [40], which used Siamese networks to learn a similarity function for generating response maps. Subsequent methods like SiamRPN++ [41], SiamR-CNN [42], SiamCAR [43], and Ocean [44] further enhanced accuracy and robustness. Recently, Vision Transformers (ViTs) [45] have introduced new possibilities for visual tracking, offering potential for streamlining and unifying tracking frameworks [45, 46, 47, 48]. For example, OStrack [48] combines feature extraction with relational modeling for efficient single-stream tracking, and HIPTrack [49] introduces a history prompting network for more precise guidance. However, most current research focuses on generic object tracking, with less attention given to specialized scenarios like camouflaged object tracking. In this work, we introduce a dataset for tracking camouflaged objects and enhance HIPTrack with multi-level features to address this challenge.
@@ II-B Visual Tracking Benchmarks
Currently, there are two main types of benchmark datasets: general benchmarks and specific benchmarks [50]. 1) General Benchmarks: Many large-scale benchmarks have been introduced for training deep learning-based trackers. For example, TrackingNet [51] aims to provide resources for training and evaluating general object tracking algorithms, containing over 30,000 sequences. GOT-10K [52] includes over 10,000 videos from the WordNet semantic hierarchy to develop the generalization capability of trackers. LaSOT [53] encompasses 1,400 long-term video sequences, with an additional 150 sequences and a new evaluation scheme. 2) Specific Benchmarks: UAV123 [54] focuses on low-altitude drone object tracking, containing 123 sequences captured by drones. CDTB [55] and PTB [56] are designed to evaluate tracking performance in RGB-D videos, where D represents depth images. TOTB [50] specializes in tracking transparent objects, collecting 225 videos from 15 categories of transparent objects. Recently, TSFMO [57] has been introduced for tracking small and fast-moving objects, while TRO [58] focuses on tracking reflected objects. Additionally, BihOT [59] contains 49 videos and 42K frames, focusing on hyperspectral camouflage. UW-COT220 [60] focuses on underwater camouflage.
@@ II-C Dealing with camouflaged Objects in Vision
Camouflaged objects blend into their surroundings due to similar textures, colors, and patterns, making them hard to distinguish. To address this, recent computer vision techniques have been developed. Fan et al. [27] introduced Camouflaged Object Detection (COD) with the COD10K dataset and the SINet framework for efficient detection. Liu et al. [28] proposed a scale-adaptive correlation filter for Camouflaged Object Segmentation (COS). Cheng et al. [22] developed a Video Camouflaged Object Detection (VCOD) framework with the MoCA-Mask dataset. Hu et al. [61] addressed visual camouflage in UAV surveillance. However, Camouflaged Object Tracking (COT), which involves consistent identification of camouflaged objects in videos, remains under-explored. To bridge this gap, we introduce the Camouflaged Object Tracking Dataset (COTD) to support algorithm development and evaluation for COT.
@@ III-A Data Collection
In the video selection process, we apply stringent criteria to ensure the high quality and practicality of the data. Specifically, videos are evaluated and selected based on three main criteria: 1) the diversity and complexity of the environment, 2) the frequency with which targets cross frame boundaries, and 3) the overall length and continuity of the videos. We prioritize videos that present additional challenges, such as a high degree of similarity between targets and backgrounds. Due to resource constraints in collecting, processing, and annotating video data, we select 200 sequences for the benchmark test. This ensures that each video is of high quality with comprehensive annotations, adhering to our available time, budget, and manpower constraints. The COTD dataset comprises 20 categories, namely: bird, bug, butterfly, cat, crab, dog, fish, frog, grasshopper, leopard, lizard, mantis, octopus, person, phasmatodea, seahorse, rabbit, shrimp, snake, and spider. To visually represent the composition of the dataset, we create a chart illustrating the number of sequences per category. Additionally, we generate a series of graphs depicting the maximum, minimum, and average number of frames for each category. These visual aids not only enhance the understanding of the dataset’s diversity but also highlight the challenges in tracking camouflaged objects across different scenarios. Fig. 3 provides a detailed view of the frame characteristics for each category, such as the range and average number of frames. These figures are crucial in demonstrating the dataset’s comprehensiveness and its potential as a robust benchmark for evaluating the performance of camouflaged object tracking algorithms. Additionally, to highlight the challenges associated with tracking camouflaged objects, we evaluate seven state-of-the-art trackers on the COTD dataset and compare them with two popular generic object tracking benchmarks. The AUC results are presented in Table I. Notably, the AUC of these trackers on the TrackingNet benchmark is more than 15.0% higher than on the COTD dataset, underscoring the significant challenges posed by tracking camouflaged objects.
@@
Figure 3: Radar chart compares min, avg, and max video frames in 20 categories. Blue, orange, green lines show distributions, highlighting differences.
@@ III-B Data Annotation
We follow the principles in [51] for sequence annotation. Each video frame is meticulously annotated by a team of expert annotators, specifically students engage in visual tracking research. Annotators draw and edit axis-aligned bounding boxes on each frame to represent the most compact boundary of the object. Visible targets are annotated as present, while those missing, out of view, or occluded are marked accordingly. To enhance the efficiency and quality of the annotations, we adopt a stringent manual annotation process. Each frame’s annotation underwent thorough review to ensure the accuracy and consistency of the bounding boxes. Annotators meticulously label each object frame-by-frame, ensuring accurate identification even within complex backgrounds. To further ensure high-quality annotations, we implement a double-check mechanism where a second annotator review the initial annotations to resolve any discrepancies. All differences are addressed through discussion and consensus. This rigorous manual annotation process and double-check mechanism guarantee the high quality of the dataset. Examples of bounding box annotations in the COTD are shown in Fig. 2.
Figure 4: Distribution of sequences with each attribute in COTD.
@@ III-C Attributes
Figure 5: The framework of HIPTrack-MLS, which inherits from HIPTrack but uses a Multi-Level Feature Sensing Encoder in the feature extraction network. The detailed structure of this encoder is shown on the right.
Following the prevalent tracking benchmark methodologies [75], [52], and [76], we meticulously annotate each video sequence with detailed attributes to enable an in-depth analysis of tracking algorithms’ performance. The proposed COTD dataset features 12 attributes, with each sequence including Background Clutter (BC) to highlight the dataset’s camouflage characteristics. In addition, we annotate the following 11 attributes: (1) Illumination Variation (IV), (2) Scale Variation (SV), (3) Deformation (DEF), (4) Motion Blur (MB), (5) Fast Motion (FM), which is marked when the object center moves by at least 20% of its own size between consecutive frames, (6) Out-of-View (OV), (7) Low Resolution (LR), marked when the object area is smaller than 900 pixels, (8) Partial Occlusion (POC), (9) Rotation (ROT), (10) Full Occlusion (FOC), and (11) Aspect Ratio Change (ARC), marked when the object’s bounding box aspect ratio falls outside the range [0.5, 2]. Fig.4 shows the distribution of these attributes on COTD. As can be seen, the most common challenge in COTD is Motion Blur. In addition, the Deformation and Partial Occlusion also present frequently in COTD. By comprehensively considering these attributes, we can conduct a thorough evaluation of tracking algorithms under various challenging conditions.